# Reinforcement Fine-Tuning of Large Language Models

### v1

@@
[Uncaptioned image]: A General-Purpose and Unified Framework for
Reinforcement Fine-Tuning of Large Language Models
Xuchen Pan, Yanxi Chen, Yushuo Chen, Yuchang Sun, Daoyuan Chen,
Wenhao Zhang, Yuexiang Xie, Yilun Huang, Yilei Zhang, Dawei Gao,
Yaliang Li, Bolin Ding, Jingren Zhou
Alibaba Group
11footnotetext: Equal contribution.22footnotetext: Corresponding author. {chenyanxi.cyx, panxuchen.pxc, yaliang.li, bolin.ding}@alibaba-inc.com
@@
Trinity-RFT is a general-purpose, flexible and scalable framework designed for reinforcement fine-tuning (RFT) of large language models. It is built with a decoupled design, consisting of (1) an RFT-core that unifies and generalizes synchronous/asynchronous, on-policy/off-policy, and online/offline modes of RFT, (2) seamless integration for agent-environment interaction with high efficiency and robustness, and (3) systematic data pipelines optimized for RFT. Trinity-RFT can be easily adapted for diverse application scenarios, and serves as a unified platform for exploring advanced reinforcement learning paradigms. This technical report outlines the vision, features, design and implementations of Trinity-RFT, accompanied by extensive examples demonstrating the utility and user-friendliness of the proposed framework.
@@
Note: Trinity-RFT is currently under active development. This technical report corresponds to commit id f17db3d (May 23, 2025) of the GitHub repository, and will be continuously updated as the codebase evolves. Comments, suggestions and contributions are welcome!
Figure 1: The design of Trinity-RFT.
1 Vision of Trinity-RFT
@@
Reinforcement learning (RL) has achieved remarkable success in the development of large language models (LLMs). Examples include aligning LLMs with human preferences via reinforcement learning from human feedback (RLHF) [18], and training long-CoT reasoning models via RL with rule-based rewards [3, 28]. However, such approaches are limited in their abilities to handle dynamic, real-world, and continuous learning.
@@
Trinity-RFT envisions a future where AI agents learn by interacting directly with environments, collecting delayed or complex reward signals, and continuously refining their behavior through RL based on the collected experiences [26]. For example, imagine an AI scientist who designs an experiment, executes it, waits for feedback (while working on other tasks concurrently), and iteratively updates itself based on true environmental rewards when the experiment is finally finished. Trinity-RFT offers a path into this future by providing various useful features.
2 Key Features
Trinity-RFT is a general-purpose, flexible, scalable and user-friendly RL framework that can be adapted for diverse experimental or real-world scenarios. The following features of Trinity-RFT make it a promising solution for realizing the aforementioned vision:
@@ 2 Key Features
  • Unified RFT modes & algorithm support. Trinity-RFT unifies and generalizes existing RFT methodologies into a flexible and configurable framework, supporting synchronous/asynchronous, on-policy/off-policy, and online/offline training, as well as hybrid modes that seamlessly combine them into a single learning process (e.g., incorporating expert trajectories to accelerate an online RL process [15, 36]). This is made possible partly by our decoupled design that allows rollout and training to be executed separately and scaled up independently on different machines, which will soon be introduced in the next section.
@@
  • Agent-environment interaction as a first-class citizen. Trinity-RFT allows delayed rewards in multi-step/time-lagged feedback loops, and handles long-tailed latencies and environment/agent failures gracefully, which ensures efficiency and robustness in real-world scenarios with complex agent-environment interaction.
  • Systematic data pipelines optimized for RFT. These include converting raw datasets to task sets for RFT, cleaning/filtering/prioritizing experiences stored in the buffer, incorporating priors and additional reward signals from various sources into the collected experiences, synthesizing data for augmenting tasks and experiences, offering user interfaces for human-in-the-loop, among others.
@@ 3.1 RFT-Core
RFT-core is the component of Trinity-RFT, highlighted at the center of Figure 1, where the core RFT process happens. Its design also exhibits a trinity, consisting of the explorer, trainer, and buffer.
@@
  • The buffer stores experiences that can be generated by the explorer or by other sources, such as human experts. It also assists with fetching training samples for the trainer, and can be integrated with advanced sampling strategies and post-processing operations.
@@
Our implementations allow the explorer and trainer to be deployed on separate machines. This decoupled design of RFT-core enables Trinity-RFT to support diverse RFT modes with great flexibility, for example:
@@
  • In a synchronous mode, the explorer and trainer get launched simultaneously, work in close coordination, and synchronize their model weights once every sync_interval training steps. Within each synchronization period, the explorer continuously generates sync_interval batches of rollout experiences and stores them in the buffer, which are then retrieved and utilized by the trainer for updating the policy model. If sync_interval $=1$, this is an on-policy RL process, whereas if sync_interval $>1$, it becomes off-policy (akin to the mode adopted in [28]) and can be accelerated by pipeline parallelism between the explorer and trainer. This mode can be activated by setting the configuration parameter mode to both.
@@
  • In a fully asynchronous mode, the explorer and trainer are mostly decoupled and can scale up independently. The explorer continuously generates rollout experiences and stores them in the buffer, while the trainer continuously samples experiences from the buffer and uses them for training the policy model. External experiences, e.g., those generated by expert models or humans, can be continuously incorporated into the buffer as well. The explorer synchronizes its model weights with the trainer once in a while, keeping the distribution of rollout experiences up to date. This mode can be activated by setting mode to explore/train and launching the explorer and trainer separately on different machines.
@@
  • Trinity-RFT supports a benchmark mode that allows the user to evaluate one or multiple checkpoints on arbitrary benchmarks, after the RFT training process has finished. To activate this mode, the user simply needs to set mode to bench and specify the paths for the evaluation datasets in the configurations. This mode can be particularly useful for experimental purposes; for example, the user might want to try out different RFT techniques or configurations quickly (with limited evaluation on hold-out data) during training, identify which RFT trials have achieved stable convergence and high rewards, and then conduct more thorough evaluations only for the checkpoints of these successful trials.
@@
  • In certain scenarios, the user would like to train the policy model without further exploration, using experiences that have already been collected and stored in the buffer. This train-only mode can be activated by setting the mode parameter to train and launching the trainer alone. Offline methods like Supervised Fine-Tuning (SFT) and Direct Preference Optimization (DPO) [20] can be regarded as special cases of such scenarios, both of which are natively supported by Trinity-RFT. For another example, consider an online RFT process that expands over a long period of time, where the explorer alone is launched during the daytime for serving human users and collecting experiences, while the trainer alone is launched at night for updating the policy model (which will be thoroughly validated and evaluated before it can be actually deployed as the rollout model for the next day).
Another benefit brought by the decoupled design is that explorers and trainers can operate across separate devices (e.g., multiple explorers on edge devices, one trainer on a high-end GPU cluster) and scale up independently. Indeed, we anticipate that for successful RL in real-world scenarios, distributed collection of massive rollout experiences could be one promising solution to resolving data scarcity.
Remark 1.
@@
We take a system-algorithm co-design perspective in the development of Trinity-RFT, and aim to unify and generalize diverse RFT methodologies in this framework, including synchronous/asynchronous, on-policy/off-policy, and online/offline modes. RFT-core provides the necessary infrastructure for achieving this goal. While the main text of this technical report focuses on the system perspective, interested readers may refer to Appendix A for further discussions on the algorithm perspective of off-policy/asynchronous RL.
@@ 3.2 Agent-Environment Interaction
Numerous challenges arise when one tries to build an RFT framework that can efficiently and robustly handle real-world interaction between the LLM-powered agent and the environment. These include long-tailed latencies, agent/environment failures, and lagged reward signals, among others.
Trinity-RFT regards agent-environment interaction as a first-class citizen and incorporates various solutions to tackle these challenges, for example:
  • Built upon vLLM [10], our implementations of the LLM inference engine in Trinity-RFT support asynchronous and concurrent generation of rollout trajectories for multiple tasks. This helps mitigate the straggler effect caused by the long-tailed latencies in rollout generation and agent-environment interaction, thereby accelerating the RFT process. Load balancing among multiple LLM inference engines within one RFT training course is also taken care of, and would be one direction for further optimizing the utilization of computational resources.
@@
To adapt Trinity-RFT to a new downstream scenario, the user mainly needs to define and register a customized workflow (by inheriting the base class Workflow or MultiTurnWorkflow) where the logic of agent-environment interaction is implemented. Detailed examples will be demonstrated in Sections 4.1 and 4.2.
@@ 3.3 Data Pipelines
The data pipelines in Trinity-RFT aim to address fundamental challenges in RFT scenarios, such as managing heterogeneous data dynamics across interaction workflows, enabling delayed reward integration, and facilitating continuous data curation. Our solutions center on three core aspects: end-to-end data transformation, active data-reward co-shaping, and human-in-the-loop curation, each corresponding to key requirements identified in our development of RFT-core (Section 3.1).

### v2

@@
Abstract
@@
Trinity-RFT is a general-purpose, unified and easy-to-use framework designed for reinforcement fine-tuning (RFT) of large language models. It is built with a modular and decoupled design, consisting of (1) an RFT-core that unifies and generalizes synchronous/asynchronous, on-policy/off-policy, and online/offline modes of RFT; (2) seamless integration for agent-environment interaction with high efficiency and robustness; and (3) systematic data pipelines optimized for RFT. Trinity-RFT can be easily adapted for diverse application scenarios, and serves as a unified platform for development and research of advanced reinforcement learning paradigms at both macroscopic and microscopic levels. This technical report outlines the vision, features, design and implementations of Trinity-RFT, accompanied by extensive examples, applications and experiments that demonstrate its functionalities and user-friendliness.
11footnotetext: Equal contribution.22footnotetext: Corresponding author. {chenyanxi.cyx,panxuchen.pxc,daoyuanchen.cdy,yaliang.li,bolin.ding}@alibaba-inc.com
@@
Note: Trinity-RFT is currently under active development. This technical report (arXiv v2 version) corresponds to commit id 63d4920 (July 14, 2025) of the GitHub repository, and will be continuously updated as the codebase evolves. Comments, suggestions and contributions are welcome!
1 Introduction
@@
Reinforcement learning (RL) has achieved remarkable success in the development of large language models (LLMs). Examples include aligning LLMs with human preferences via reinforcement learning from human feedback (RLHF) [23], and training long-CoT reasoning models via RL with rule-based or verifiable rewards (RLVR) [5, 34]. However, such approaches are limited in their abilities to handle dynamic, agentic and continuous learning in the real world.
@@
We envision a future where AI agents learn by interacting directly with environments, collecting lagged or complex reward signals, and continuously refining their behavior through RL based on the collected experiences [31]. For example, imagine an AI scientist who designs an experiment, executes it, waits for feedback (while working on other tasks concurrently), and iteratively updates itself based on true environmental rewards and feedback when the experiment is finally finished.
This vision motivates us to develop Trinity-RFT, a reinforcement fine-tuning (RFT) framework that aims to offer a path into this future. The modular, decoupled and trinity design of Trinity-RFT illustrated in Figure 1, along with its various features, makes it a promising solution for realizing such a vision.
Figure 1: The high-level design of Trinity-RFT.
1.1 Key Features of Trinity-RFT
Trinity-RFT is a general-purpose, unified, scalable and user-friendly RL framework that can be easily adapted for diverse experimental or real-world scenarios. It integrates both macroscopic and microscopic RL methodologies in one place; roughly speaking, the former deals with natural language and plain text, while the latter handles torch.Tensor (such as token probabilities, gradients, and model weights of LLMs) [^1] Many prior RL works for games/control/LLMs focus mainly on the microscopic aspect, e.g., designing policy loss functions or optimization techniques for updating the policy model. On the other hand, pre-trained LLMs, as generative models with rich prior knowledge of natural language and the world, open up numerous opportunities at the macroscopic level, e.g., experience synthesis by reflection or reasoning with environmental feedback [4], leveraging existing text processing methods like deduplication and quality filtering [2], among others.. The key features of Trinity-RFT are presented below, which will be further elaborated in Section 2 and exemplified in Section 3.
An RFT-core that unifies and generalizes diverse RL modes.
@@ 1.1 Key Features of Trinity-RFT
Trinity-RFT implements diverse RL methodologies in a unified manner, supporting synchronous/asynchronous, on-policy/off-policy, and online/offline training. These RL modes can be flexibly generalized, e.g., a hybrid mode that incorporates expert trajectories to accelerate an online RL process [20, 45]. This unification is made possible partly by our decoupled design (which will soon be introduced in Section 2.1) that allows rollout and training to be executed separately and scaled up independently on different devices, while having access to the same stand-alone experience buffer. The efficacy of various RL modes has been validated empirically by our experiments in Section 3.3, which particularly highlight the efficiency gains by off-policy or asynchronous methods.
Agent-environment interaction as a first-class citizen.
@@
Trinity-RFT allows delayed rewards and environmental feedback in multi-step/time-lagged feedback loop, handles long-tailed latencies and the straggler effect via asynchronous and streaming LLM inference, and deals with environment/agent failures gracefully via dedicated timeout/retry/skip mechanisms. These together ensure efficiency and robustness of continuous agent-environment interaction in complex real-world scenarios.
Systematic data pipelines optimized for RFT.
Figure 2 illustrates the high-level design of data pipelines in Trinity-RFT, which regard rollout tasks and experiences as dynamic assets to be actively managed throughout the RFT lifecycle. Trinity-RFT empowers users to: (1) curate tasks for curriculum learning, e.g., by prioritizing easier tasks at the beginning of training to stabilize and accelerate the learning process; (2) actively manipulate experience by cleaning, filtering, or synthesizing new experiences, such as repairing failed trajectories or amplifying successful ones; (3) perform online reward shaping by augmenting sparse environmental rewards with dense, computed signals, such as quality or diversity scores; (4) customize interfaces for human-in-the-loop curation and utilize an agentic paradigm for RFT data processing that translates high-level natural language commands (e.g., “improve response diversity and safety for coding scenarios”) into complex data pipelines, powered by established community tools like Data-Juicer [2]. For instance, Section 3.4 presents experiments that demonstrate the efficacy of task prioritization and reward shaping empowered by Trinity-RFT.
Figure 2: The high-level design of data pipelines in Trinity-RFT.
User-friendliness as a top priority.
For development and research, the modular and decoupled design of Trinity-RFT allows the user to develop new RFT methodologies by adding one or a few small, plug-and-play classes (modified from built-in templates) that implement the essential functionalities of interest, with minimal code duplication or intrusive changes to the codebase. An example can be found in Section 3.2, which shows that three compact python classes (with around 200 lines of code in total) suffice for implementing a hybrid RL process that leverages samples from multiple data sources and updates the policy model with a customized loss function. For applications, the user can adapt Trinity-RFT to a new scenario by simply implementing a single Workflow class that specifies the logic of agent-environment interaction, as will be exemplified in Section 3.1. To further enhance usability, Trinity-RFT incorporates various graphical user interfaces to support low-code usage and development, enhance transparency of the RFT process, and facilitate easy monitoring and tracking.
1.2 Related Works
There exist numerous open-source RLHF frameworks, such as veRL [29], OpenRLHF [12], TRL [39], ChatLearn [1], Asynchronous RLHF [22], among others. Some of them have been further adapted for training long-CoT reasoning models or for agentic RL more recently.
Concurrent to Trinity-RFT, some recent works on LLM reinforcement learning also advocate a decoupled and/or asynchronous design; examples include StreamRL [49], MiMo [43], AReaL [8], ROLL [36], LlamaRL  [42], Magistral [17], AsyncFlow [11], among others.
Complementary to this huge number of related works, Trinity-RFT provides the community with a new solution that is powerful, easy-to-use, and unique in certain aspects. In a nutshell, Trinity-RFT aims to be general-purpose and applicable to diverse application scenarios, while unifying various RFT modes, RFT methodologies at macroscopic and microscopic levels, and RFT-core/agent-environment interaction/data pipelines. Such a system-engineering perspective makes Trinity-RFT particularly useful for handling the whole RFT pipeline in one place. We also hope that some specific features of Trinity-RFT, such as data persistence in the experience buffer, and distributed deployment of multiple independent explorers, will open up new opportunities for LLM reinforcement fine-tuning.
@@ 2.1 RFT-Core
RFT-core is the component of Trinity-RFT, highlighted at the center of Figure 1, where the core RFT process happens. Its design also exhibits a trinity, consisting of the explorer, buffer, and trainer.
@@
  • The buffer stores experiences that can be generated by the explorer or by other sources, such as human experts. It can be realized in various forms, such as a non-persistent ray.Queue or a persistent SQLite database. It also assists with fetching training samples for the trainer, and can be integrated with advanced sampling strategies and post-processing operations.
@@
Our implementations allow the explorer and trainer to be deployed on separate machines and act independently. They are only connected via (1) access to the same experience buffer with a customizable sampling strategy, and (2) model weight synchronization by a customizable schedule. See Figure 3 for an illustration. This decoupled design of RFT-core offers support for diverse RFT modes with great flexibility.
Figure 3: The architecture of RFT-core in Trinity-RFT.
2.1.1 Unified Support for Diverse RFT Modes
We present the RFT modes supported by Trinity-RFT, some of which are demonstrated in Figure 4.
Figure 4: A visualization of diverse RFT modes supported by Trinity-RFT, including: (a) synchronous mode, with sync_interval=2; (b) one-step off-policy mode, with sync_interval=1 and sync_offset=1; (c) fully asynchronous mode, with sync_interval=2; (d) multi-explorer asynchronous mode, with sync_interval=2. The buffer supports, in principle, arbitrary management and sampling strategies for experiences.
Synchronous mode.
@@ 2.1.1 Unified Support for Diverse RFT Modes
In the synchronous mode shown in Figure 4 (a), the explorer and trainer get launched simultaneously, work in close coordination, and synchronize their model weights once every sync_interval training steps. Within each synchronization period, the explorer continuously generates sync_interval batches of rollout experiences and stores them in the buffer, which are then retrieved and utilized by the trainer for updating the policy model. If sync_interval=1, this is a strictly on-policy RL process, whereas if sync_interval>1, it becomes off-policy (akin to the mode adopted in [34]) and can be accelerated by pipeline parallelism between the explorer and trainer. This mode can be activated by setting the configuration parameter mode=both.
One-step off-policy mode.
This mode, demonstrated in Figure 4 (b), closely resembles the synchronous mode, except for an offset of one batch between the explorer and trainer. This allows the trainer to sample experiences from the buffer immediately after model weight synchronization, thereby streamlining the execution of explorer and trainer with smaller pipeline bubbles, at the cost of slight off-policyness. The visualization in Figure 4 (b) corresponds to configuration parameters sync_interval=1 and sync_offset=1, both of which can take more general values in Trinity-RFT.
Asynchronous mode.
@@
In the fully asynchronous mode shown in Figure 4 (c), the explorer and trainer act almost independently. The explorer continuously generates rollout experiences and stores them in the buffer, while the trainer continuously samples experiences from the buffer and uses them for training the policy model. External experiences, e.g., those generated by expert models or humans, can be continuously incorporated into the buffer as well. The explorer and trainer independently load or save model weights from the checkpoint directory every sync_interval steps, keeping the distribution of rollout experiences up to date. This mode can be activated by setting mode=explore/train and launching the explorer and trainer separately on different GPUs.
Multi-explorer asynchronous mode.
One benefit brought by the decoupled design is that explorers and trainers can scale up independently on separate devices. As a proof-of-concept, Trinity-RFT offers support for a multi-explorer asynchronous mode, as demonstrated in Figure 4 (d), where multiple explores send the generated rollout experiences to the same buffer. Scaling up the number of independent and distributed explorers can be particularly useful for resolving data scarcity and speeding up the generation of experiences in real-world scenarios where rollout trajectories have to be sampled via interaction with the physical world, or in an environment with sparse and lagged feedback. Another by-product of this multi-explorer mode is 24-hour non-interrupted service for real-world online serving situations: since the explorers can pause and update model weights at different moments, it can be guaranteed that there is always one explorer ready to serve an incoming request immediately whenever it arrives. This is in contrast to a single-explorer mode, where online service has to be paused when the explorer is updating its model weights.
Benchmark mode.
@@
Trinity-RFT supports a benchmark mode that allows the user to evaluate one or multiple checkpoints on arbitrary benchmarks, after the RFT training process has finished. To activate this mode, the user simply needs to set mode=bench and specify the paths for the evaluation datasets in the configurations. This mode can be particularly useful for experimental purposes; for example, the user might want to try out different RFT techniques or configurations quickly (with limited evaluation on hold-out data) during training, identify which RFT trials have achieved stable convergence and high rewards, and then conduct more thorough evaluations only for the checkpoints of these successful trials.
Train-only mode.
@@
In certain scenarios, the user would like to train the policy model without further exploration, using experiences that have already been collected and stored in the buffer. This train-only mode can be activated by setting the configuration parameter mode $=$ train and launching the trainer alone. Offline methods like Supervised Fine-Tuning (SFT) and Direct Preference Optimization (DPO) [24] can be regarded as special cases of such scenarios, both of which are natively supported by Trinity-RFT. For another example, consider an online RFT process that expands over a long period, where the explorer alone is launched during the daytime for serving human users and collecting experiences, while the trainer alone is launched at night for updating the policy model (which will be thoroughly validated and evaluated before it can be actually deployed as the rollout model for the next day).
Discussions.
@@
We conclude this subsection with two remarks. (1) Given the unified implementation of various RFT modes, it is easy to design and implement a hybrid mode with Trinity-RFT that combines multiple modes into a single learning process. One example is learning with both online rollout data and offline-collected expert data, via jointly optimizing two loss terms corresponding to these two data sources. Section 3.2 illustrates how to implement this conveniently in Trinity-RFT. (2) We take a system-algorithm co-design perspective in the development of Trinity-RFT, aiming to unify and generalize diverse RFT methodologies in this framework. RFT-core provides the necessary infrastructure for achieving this goal. While the main text of this technical report focuses on the system perspective, interested readers may refer to Appendix A for further discussions on the algorithm perspective of off-policy/asynchronous RL.
2.1.2 Implementations of RFT-Core
We present some implementation details of RFT-core in the following.
Inference and training engines.
The current version of Trinity-RFT leverages vLLM [14] as the inference engine for the explorer, which offers features including paged attention, continuous batching [48], asynchronous and concurrent inference for multiple rollout trajectories, among others. Trinity-RFT also leverages verl [29] as the training engine for the trainer, which gracefully handles model placement (for the policy, critic and reference models) and incorporates various performance optimizations for training (such as dynamic batching, management of padding and unpadding, etc.). Trinity-RFT stands on the shoulders of these excellent open-source projects, and will continue to benefit from their future development.
Experience buffer.
Trinity-RFT supports multiple types of experience buffers, ranging from a non-persistent ray.Queue to persistent SQLite or Redis databases. While using a basic first-in-first-out queue is the most straightforward approach, data persistence with a database opens up many new opportunities (e.g., advanced sampling strategies), as discussed throughout this report. Trinity-RFT has provided dedicated read/write control to prevent any conflict in accessing the buffer.
Model weight synchronization.
Trinity-RFT supports model weight synchronization between the explorer and trainer by NCCL [21], or by checkpoint saving and loading. The former is faster (when available), while the latter is generally more flexible and widely applicable, especially for asynchronous RFT modes.
@@ 2.2 Agent-Environment Interaction
To adapt Trinity-RFT to a new downstream scenario, the user mainly needs to define and register a customized workflow (by inheriting the base class Workflow or MultiTurnWorkflow) where the logic of agent-environment interaction for this particular scenario is implemented. Advanced methods for experience synthesis with environmental feedback [4] can be implemented in the same way as well. See Section 3.1 for detailed examples. The workflow will then be executed by workflow runners within the explorer for generating experiences, as shown in Figure 3.
Numerous challenges arise when one tries to build an RFT framework that can efficiently and robustly handle real-world interaction between the LLM-powered agent and the environment. These include long-tailed latencies, agent/environment failures, and lagged reward signals, among others. Trinity-RFT regards agent-environment interaction as a first-class citizen and incorporates various solutions to tackle these challenges, for example:
  • The workflow runners in Trinity-RFT support asynchronous and streaming generation of rollout trajectories for multiple tasks. This helps mitigate the straggler effect caused by the long-tailed latencies in rollout generation and agent-environment interaction, thereby accelerating the RFT process. Load balancing among multiple LLM inference engines within one RFT training course is also taken care of, and would be one direction for further optimizing the utilization of computational resources.
@@
  • For multi-turn conversations and ReAct-style workflows [47], Trinity-RFT supports concatenating multiple rounds of agent-environment interaction compactly into a single sequence, with proper masking that indicates which tokens need to be incorporated into the training objective of RL algorithms. This avoids unnecessary recomputation and thus improves training efficiency, compared to a vanilla approach that represents a $K$-turn rollout trajectory with $K$ separate samples.
  • As another performance optimization, the implementation of Trinity-RFT allows resetting the environment in a workflow, rather than re-initializing it every time. This is especially useful for scenarios where setting up the environment is costly.
@@ 2.3 Data Pipelines
The data pipelines in Trinity-RFT aim to address fundamental challenges in RFT scenarios, such as managing heterogeneous data dynamics across interaction workflows, enabling delayed reward integration, and facilitating continuous data curation. Our solutions center on four core aspects: end-to-end data transformation, task curation, active experience shaping, and human-in-the-loop curation, each corresponding to key requirements identified in our development of RFT-core (Section 2.1).